Find the area bounded by the given curves, axes and/or lines

y2 = x , y = x
1
Expert's answer
2021-10-18T08:25:56-0400





Area enclosed =ย "\int^1_0 \sqrt{x}. dx-\int^1_0 x.dx\\"

"=[\frac{x^{\frac{3}{2}}}{\frac{3}{2}}-\frac{x^2}{2}]^1_0\\
=[\frac{2}{3}-\frac{1}{2}]-[0-0]\\
=\frac{1}{6}\\"